OpenStudy (anonymous):

is the answer to "7x-1=37 +8" y=7/3x - 3?

OpenStudy (anonymous):

well from the looks of this question you probably just have to find what x is.

OpenStudy (anonymous):

its supposed to be in y intercept form

OpenStudy (anonymous):

7x-1=37 +8 7x-1=45 7x=46 x=6.57

OpenStudy (anonymous):

ok can you write out the whole problem with directions please